OpenStudy (anonymous):

the perimiter of a rectangle is 112 inches.the length exceeds the width by 16 inches. find length and width

jhonyy9 (jhonyy9):

112=2l+2w l=w+16

OpenStudy (anonymous):

I do know nabout the 2l and 2w i just cant figure it out ...flustered

jhonyy9 (jhonyy9):

112=2(w+16)+2w 112=2w+32+2w 80=4w w=20 in l=20+16=36 112=36+36+20+20